# Relation between equilibrium quantum phase transitions and dynamical quantum phase transitions in two-band systems

### v1

@@ Abstract
The dynamical quantum phase transition (DQPT) is an important concept in non-equilibrium critical phenomena, however, its relation to the equilibrium quantum phase transition (EQPT) remains obscure. Substantial evidences have suggested that quenching across the underlying equilibrium phase boundary is neither a sufficient nor a necessary condition for the existence of the DQPT. In this work, we give a necessary and sufficient condition for the occurrence of DQPTs in two-band systems by introducing the quench fidelity, which is defined as the fidelity between the ground state wavefunctions of the pre-quench and post-quench Hamiltonian, and elaborate it by taking one-dimensional anisotropic XY model as an example. The relation between EQPTs and DQPTs is analyzed in detail in terms of the quench fidelity.
@@ I Introduction
The equilibrium quantum phase transition (EQPT) is caused by the quantum fluctuations and controlled by the parameters of the Hamiltonian of the system Sachdev . The existence of the EQPT reveals that quantum systems with parameters in different equilibrium quantum phases have distinct ground-state properties Zanardi2006 ; You2007 . The fidelity, which is the module of the overlap of two ground states of Hamiltonians with different parameter values, is widely used to describe the non-analytic changes of the properties of ground states around EQPT points Zanardi2006 ; You2007 ; Zanardi2007 ; Buonsante2007 ; Cozzini2007-1 ; Cozzini2007-2 ; YanfMF2007 ; ZhouHQ2008 ; ChenS2008 ; ZRams2011 ; Zeng2024 ; Dutta ; SunG2015 ; SunG . Using $\gamma=(\gamma_{1},\gamma_{2},\ldots)$ to denote phase-driving parameters of the quantum system, the fidelity can be formulated as
$\mathcal{F}(\gamma,\tilde{\gamma})=\left|\left\langle\psi^{0}(\gamma)\right|\left.\psi^{0}(\tilde{\gamma})\right\rangle\right|,$ (1)
where $\left|\psi^{0}(\gamma)\right\rangle$ denotes the ground state of the Hamiltonian $H(\gamma)$ and $\tilde{\gamma}$ represents parameters with values different from $\gamma$. A previous work has unveiled that $\mathcal{F}(\gamma,\tilde{\gamma})$ has (has not) exact zeros in the thermodynamical limit for $\gamma$ and $\tilde{\gamma}$ are in different (the same) phases Zeng2024 . To eliminate the effect of the Anderson orthogonality catastrophe Anderson1967 (OC) in the thermodynamical limit, it is natural to introduce the decay rate function given by
$\alpha(\gamma,\tilde{\gamma})=-\frac{1}{L}\ln\mathcal{F}(\gamma,\tilde{\gamma}),$ (2)
where $L$ is the size of the system. Fixing the value of $\gamma$ and changing the value of $\tilde{\gamma}$, $\alpha(\tilde{\gamma})$ will exhibit non-analytical behavior in the thermodynamical limit at EQPT points. Thus the decay rate function $\alpha(\tilde{\gamma})$ is a valid tool to probe the EQPT.
Since the dynamical quantum phase transition (DQPT) was proposed more than a decade ago Heyl2013PRL , it has been widely recognized that the DQPT is closely related to the EQPT Karrasch2013PRB ; Kriel2014PRB ; Heyl2015RPL ; Dora2015PRB ; Schmitt2015PRB ; HuangZ2016PRB ; Zvyagin2016LTP ; Bhattacharya2017PRB ; Bhattacharjee2018PRB ; Heyl2018RPP ; Cheraghi2023NJP ; Corps ; ZhouBZ2019 . The DQPT is a phenomenon in non-equilibrium systems which indicates that the time-evolution state evolves into an orthogonal state of the initial state. As an analogue of the fidelity in EQPT theory, the Loschmidt echo (LE) is a key concept in DQPT theory. For an initial state $|\psi_{i}\rangle$ which is the ground state of the initial Hamiltonian $H_{i}$ and its time evolution state $|\psi(t)\rangle=e^{-iH_{f}t}|\psi_{i}\rangle$ which is driven by a post-quench Hamiltonian $H_{f}$, the LE is defined as
@@
which represents the return probability of the time-evolution state to the initial state. Similar to the fidelity at EQPT points, the LE has exact zeros at critical times when DQPTs happen ZhouBZ2021 ; ZengYM . Similarly, to remove the effect of system size properly, the rate function of the LE is introduced as
@@
where $L$ is the size of the system. Evidently, the exact zero of the LE $\mathcal{L}(t)$ can give rise to the singularity of the rate function $\lambda(t)$. Therefore, DQPTs can be characterized by non-analytic behaviors of the rate function $\lambda(t)$ of the LE in the real-time evolution.
In previous works Heyl2013PRL ; Karrasch2013PRB ; Kriel2014PRB ; Bhattacharjee2018PRB , it is commonly believed that the condition for the occurrence of DQPTs is consistent with EQPTs, i.e., if the phase-driving parameters are suddenly quenched across the underlying equilibrium phase boundary at the initial time $t=0$, DQPTs will arise at certain critical times. In noninteracting topological systems, it is clear that DQPTs have to occur if the topological number changes under the quench in 1D systems Dora2015PRB ; YangC , while the occurrence of DQPTs are related to the change in the absolute value of the Chern number in 2D systems HuangZ2016PRB . These types of DQPTs are robust and are therefore termed topologically protected. However, it has been discovered that certain particular models do have DQPTs even for pre-qunch and post-quench parameters in the same equilibrium phase, which are termed accidental DQPTs Dora2015PRB ; Liska ; Heyl2018RPP ; WongCY2024PRB . In fact, quenching across the underlying equilibrium phase boundary is neither a sufficient nor a necessary condition for the existence of the DQPT Vajna2014PRB ; Andraschko2014PRB ; Sharma2015PRB ; Divakaran2016PRE ; Obuchi2017PRB ; Halimeh2017PRB ; Zauner-Stauber2017PRE ; Jafari2019SR ; Sadrzadeh2021PRB ; Stumper2022PRR ; Rossi2022PRB . It means that DQPT may not happen even quenching across the underlying EQPT point. So here comes the question: is there an intrinsic relation between EQPTs and DQPTs?
In this paper, we expose the relation between EQPTs and DQPTs in two-band systems from a new perspective by introducing a concept of the quench fidelity. In terms of the quench fidelity, we give a universal necessary and sufficient condition for the occurrence of DQPTs. Then, we demonstrate our theory by taking the XY model as a concrete example and calculating it in detail. Through analysis of the distribution of values of each $k$-mode of the quench fidelity, we also propose a sufficient condition for the occurrence of DQPTs. At last, we discuss the relationship between quenching across the equilibrium phase transition point and the existence of DQPTs.
@@ II The quench fidelity
In order to discuss DQPTs and EQPTs in the same framework, we shall introduce the definition of the quench fidelity and reveal its role in connecting EQPTs and DQPTs.
@@
where $H_{k}(\gamma)$ is the Hamiltonian of $k$-mode with momentum $k$; $\boldsymbol{\mathbf{\sigma}}=(\sigma_{x},\sigma_{y},\sigma_{z})$ are Pauli matrices; $\mathbf{d}_{k}(\gamma)=(d_{x,k}(\gamma),d_{y,k}(\gamma),d_{z,k}(\gamma))$ and $d_{0,k}(\gamma)$ are corresponding vector components of $H_{k}(\gamma)$; and $\mathbb{I}$ denotes the unit matrix. Choosing $H_{i}=H(\gamma_{i})$ as the initial Hamiltonian, where $\gamma_{i}$ is the pre-quench parameter. At time $t=0$, suddenly quenching the value of $\gamma$ from $\gamma_{i}$ to $\gamma_{f}$, i.e., $\gamma=\gamma_{i}\Theta(-t)+\gamma_{f}\Theta(t)$. Then the LE of the system which is evolved under the post-quench Hamiltonian $H_{f}=H(\gamma_{f})$ can be represented as
@@
where $H_{f,k}$ is the Hamiltonian of $k$-mode with parameter $\gamma_{f}$, and $|\psi_{i,k}^{0}\rangle$ is the ground state of the pre-quench Hamiltonian $H_{i,k}$ of $k$-mode with parameter $\gamma_{i}$. Then we have the $k$-mode of the LE:
@@
The occurrence of the DQPT requires at least one $k_{c}$-mode of the LE equals to 0, i.e., $\mathcal{L}_{k_{c}}(t_{c,n})=0$, where $n$ is a positive integer and $t_{c,n}=(2n-1)\pi/(2\left|\mathbf{d}_{k_{c}}(\gamma_{f})\right|)$ are critical times when DQPTs happen ZengYM . Since $\sin^{2}(\left|\mathbf{d}_{k_{c}}(\gamma_{f})\right|t_{c,n})$ always equals to 1, it is easy to get $\mathcal{L}_{k_{c}}(t_{c,n})=(\hat{\mathbf{d}}_{k_{c}}(\gamma_{i})\cdot\mathbf{\hat{d}}_{k_{c}}(\gamma_{f}))^{2}$. Evidently, $\mathcal{L}_{k_{c}}(t_{c,n})=0$ is fulfilled if and only if $\hat{\mathbf{d}}_{k_{c}}(\gamma_{i})\cdot\mathbf{\hat{d}}_{k_{c}}(\gamma_{f})=0$, .i.e., $\hat{\mathbf{d}}_{k_{c}}(\gamma_{i})$ and $\mathbf{\hat{d}}_{k_{c}}(\gamma_{f})$ are perpendicular on the Bloch sphere Dora2015PRB . Whereas what we care about is whether such a $k_{c}$ exists for various values of $\gamma_{i}$ and $\gamma_{f}$, we take no account of values of critical times $t_{c,n}$ when $\mathcal{L}_{k_{c}}$ equals to zero, and investigate properties of ¯$\bar{\mathcal{L}}_{k}=\mathcal{L}_{k}(\pi/(2\left|\mathbf{d}_{k}(\gamma_{f})\right|))$ which is the minimum of $\mathcal{L}_{k}(t)$ in time instead of $\mathcal{L}_{k}(t)$. According to Eq. (7), it is easy to get
¯$\bar{\mathcal{L}}_{k}=(\hat{\mathbf{d}}_{k}(\gamma_{i})\cdot\mathbf{\hat{d}}_{k}(\gamma_{f}))^{2},$ (8)
which is time independent and equals to zero only for the $k_{c}$-mode, i.e., ¯$\bar{\mathcal{L}}_{k_{c}}=0$.
@@
where $|\psi_{i,k}^{0}\rangle$ and $|\psi_{f,k}^{0}\rangle$ are ground states of pre-quench Hamiltonian $H_{i,k}$ and post-quench Hamiltonian $H_{f,k}$, respectively. Comparing Eq. (9) with Eq. (1), the difference is that the two parameters of the quench fidelity are fixed to pre-quench parameter and post-quench parameter, respectively. Then we have the $k$-mode of the quench fidelity:
@@
For the case $\gamma_{i}$ and $\gamma_{f}$ are in different quilibrium phases, $\mathcal{F}^{\text{q}}$ has at least one $k_{0}$-mode that fulfills $\mathcal{F}^{\text{q}}_{k_{0}}=0$, thus there exists at least one $k_{0}$-mode that fulfills $\hat{\mathbf{d}}_{k_{0}}(\gamma_{i})\cdot\mathbf{\hat{d}}_{k_{0}}(\gamma_{f})=-1$, .i.e., $\hat{\mathbf{d}}_{k_{0}}(\gamma_{i})$ and $\hat{\mathbf{d}}_{k_{0}}(\gamma_{f})$ are antiparallel on the Bloch sphere.
Comparing Eq. (10) whit Eq. (8), it is straightforward to get
¯$\bar{\mathcal{L}}_{k}=\left[2(\mathcal{F}^{\text{q}}_{k})^{2}-1\right]^{2}.$ (11)
Eq. (11) implies that the $k_{c}$-mode of the LE which satisfies ¯$\bar{\mathcal{L}}_{k_{c}}=0$ corresponds to the $k_{c}$-mode of $\mathcal{F}^{\text{q}}$ which satisfies $\mathcal{F}^{\text{q}}_{k_{c}}=\sqrt{2}/2$. Therefore, a necessary and sufficient condition for the occurrence of DQPTs is the quench fidelity of the pre-quench system and the post-quench system has at least a $k_{c}$-mode which satisfies $\mathcal{F}^{\text{q}}_{k_{c}}=\sqrt{2}/2$. For a two-band system, if $d_{0,k_{c}}(\gamma_{f})=0$, then
$H_{k_{c}}(\gamma_{f})\left|\psi_{k_{c}}^{0}(\gamma_{i})\right\rangle\propto\left|\psi_{k_{c}}^{1}(\gamma_{i})\right\rangle$
can be derived from $\mathcal{F}_{k_{c}}=\sqrt{2}/2$, where $\left|\psi_{k_{c}}^{1}(\gamma_{i})\right\rangle$ is the first excited state of $H_{k_{c}}(\gamma_{i})$. It implies $\left\langle\psi_{k_{c}}^{0}(\gamma_{i})\right|H_{k_{c}}(\gamma_{f})\left|\psi_{k_{c}}^{0}(\gamma_{i})\right\rangle=0$. And further $\left\langle\psi_{k_{c}}^{0}(\gamma_{i})\right|H_{k_{c}}^{2}(\gamma_{f})\left|\psi_{k_{c}}^{0}(\gamma_{i})\right\rangle=\left|\mathbf{d}_{k_{c}}(\gamma_{f})\right|^{2}$ can be derived. If $d_{0,k_{c}}(\gamma_{i})$ is also equal to 0, then $H_{k_{c}}(\gamma_{i})$ anti-commutes with $H_{k_{c}}(\gamma_{f})$, i.e.,
$\{H_{k_{c}}(\gamma_{i}),H_{k_{c}}(\gamma_{f})\}=0.$
The detailed derivation of these formulas is given in Appendix A.

### v2

@@ Abstract
The dynamical quantum phase transition (DQPT) is an important concept in nonequilibrium critical phenomena; however, its relation to the equilibrium quantum phase transition (EQPT) remains obscure. Substantial evidence has suggested that quenching across the underlying equilibrium phase boundary is neither a sufficient nor a necessary condition for the existence of the DQPT. In this work, we give a necessary and sufficient condition for the occurrence of DQPTs in two-band systems by introducing the quench fidelity, which is defined as the fidelity between the ground-state wave functions of the prequench and postquench Hamiltonian, and elaborate it by taking the one-dimensional anisotropic XY model as an example. The relation between EQPTs and DQPTs is analyzed in detail in terms of the quench fidelity.
@@ I Introduction
The equilibrium quantum phase transition (EQPT) is caused by the quantum fluctuations and controlled by the parameters of the Hamiltonian of the system Sachdev . The existence of the EQPT reveals that quantum systems with parameters in different equilibrium quantum phases have distinct ground-state properties Zanardi2006 ; You2007 . The fidelity, which is the module of the overlap of two ground states of Hamiltonians with different parameter values, is widely used to describe the nonanalytic changes of the properties of ground states around EQPT points Zanardi2006 ; You2007 ; Zanardi2007 ; Buonsante2007 ; Cozzini2007-1 ; Cozzini2007-2 ; YanfMF2007 ; ZhouHQ2008 ; ChenS2008 ; ZRams2011 ; Zeng2024 ; Dutta ; SunG2015 ; SunG . Using $\gamma=(\gamma_{1},\gamma_{2},\ldots)$ to denote phase-driving parameters of the quantum system, the fidelity can be formulated as
[…]
Since the dynamical quantum phase transition (DQPT) was proposed more than a decade ago Heyl2013PRL , it has been widely recognized that the DQPT is closely related to the EQPT Karrasch2013PRB ; Kriel2014PRB ; Heyl2015RPL ; Dora2015PRB ; Schmitt2015PRB ; HuangZ2016PRB ; Zvyagin2016LTP ; Bhattacharya2017PRB ; Bhattacharjee2018PRB ; Heyl2018RPP ; Cheraghi2023NJP ; Corps ; ZhouBZ2019 . The DQPT is a phenomenon in nonequilibrium systems which indicates that the time-evolution state evolves into an orthogonal state of the initial state. As an analog of the fidelity in EQPT theory, the Loschmidt echo (LE) is a key concept in DQPT theory. For an initial state $|\psi_{i}\rangle$ which is the ground state of the initial Hamiltonian $H_{i}$ and its time evolution state $|\psi(t)\rangle=e^{-iH_{f}t}|\psi_{i}\rangle$ which is driven by a postquench Hamiltonian $H_{f}$, the LE is defined as
@@
which represents the return probability of the time-evolution state to the initial state. Similar to the fidelity across EQPT points, the LE has exact zeros at critical times when DQPTs happen ZhouBZ2021 ; ZengYM . Similarly, to remove the effect of system size properly, the rate function of the LE is introduced as
@@
where $L$ is the size of the system. Evidently, the exact zero of the LE $\mathcal{L}(t)$ can give rise to the singularity of the rate function $\lambda(t)$. Therefore, DQPTs can be characterized by nonanalytic behaviors of the rate function $\lambda(t)$ of the LE in the real-time evolution.
In early works Heyl2013PRL ; Karrasch2013PRB ; Kriel2014PRB ; Bhattacharjee2018PRB , it was commonly believed that the condition for the occurrence of DQPTs is consistent with EQPTs; i.e., if the phase-driving parameters are suddenly quenched across the underlying equilibrium phase boundary at the initial time $t=0$, DQPTs will arise at certain critical times. In noninteracting topological systems, it is clear that DQPTs have to occur if the topological number changes under the quench in 1D systems Dora2015PRB ; YangC , while the occurrence of DQPTs is related to the change in the absolute value of the Chern number in 2D systems HuangZ2016PRB . These types of DQPTs are robust and are therefore termed topologically protected. However, it has been discovered that certain particular models do have DQPTs for quenches within the same equilibrium phase, and even for very weak quenches, which are termed accidental DQPTs Dora2015PRB ; Liska ; Flaschner ; Heyl2018RPP ; WongCY2024PRB and anomalous DQPTs Homrighausen2017 ; Halimeh2020 . In fact, quenching across the underlying equilibrium phase boundary is neither a sufficient nor a necessary condition for the existence of the DQPT Vajna2014PRB ; Andraschko2014PRB ; Sharma2015PRB ; Divakaran2016PRE ; Obuchi2017PRB ; Halimeh2017PRB ; Zauner-Stauber2017PRE ; Jafari2019SR ; Sadrzadeh2021PRB ; Stumper2022PRR ; Rossi2022PRB . This means that a DQPT may not happen even quenching across the underlying EQPT point. The intrinsic relation between EQPTs and DQPTs is an intriguing question. Although some works Lang2018PRB ; Uhrich2020PRB have discussed this issue, a theoretical study to uncover clearly their relation is necessary.
In this paper, we expose the relation between EQPTs and DQPTs in a general two-band system from a new perspective by introducing a concept of the quench fidelity. In terms of the quench fidelity, we give a universal necessary and sufficient condition for the occurrence of DQPTs. Then, we demonstrate our theory by taking the XY model as a concrete example and calculating it in detail. Through analysis of the distribution of values of each $k$-mode of the quench fidelity, we also propose a sufficient condition for the occurrence of DQPTs. Finally, we discuss the relationship between quenching across the equilibrium phase transition point and the existence of DQPTs.
@@ II The quench fidelity
In order to discuss DQPTs and EQPTs in the same framework, we will introduce the definition of the quench fidelity and reveal its role in connecting EQPTs and DQPTs.
@@
where $H_{k}(\gamma)$ is the Hamiltonian of $k$-mode with momentum $k$, $\boldsymbol{\mathbf{\sigma}}=(\sigma_{x},\sigma_{y},\sigma_{z})$ are Pauli matrices, $\mathbf{d}_{k}(\gamma)=(d_{x,k}(\gamma),d_{y,k}(\gamma),d_{z,k}(\gamma))$ and $d_{0,k}(\gamma)$ are corresponding vector components of $H_{k}(\gamma)$, and $\mathbb{I}$ denotes the unit matrix. We choose $H_{i}=H(\gamma_{i})$ as the initial Hamiltonian, where $\gamma_{i}$ is the prequench parameter. At time $t=0$, we suddenly quench the value of $\gamma$ from $\gamma_{i}$ to $\gamma_{f}$, i.e., $\gamma=\gamma_{i}\Theta(-t)+\gamma_{f}\Theta(t)$. Then the LE of the system which is evolved under the postquench Hamiltonian $H_{f}=H(\gamma_{f})$ can be represented as
@@
where $H_{f,k}$ is the Hamiltonian of the $k$-mode with parameter $\gamma_{f}$, and $|\psi_{i,k}^{0}\rangle$ is the ground state of the prequench Hamiltonian $H_{i,k}$ of the $k$-mode with parameter $\gamma_{i}$. Then we have the $k$-mode of the LE:
@@
The occurrence of the DQPT requires at least one $k_{c}$-mode of the LE equal to 0, i.e., $\mathcal{L}_{k_{c}}(t_{c,n})=0$, where $n$ is a positive integer and $t_{c,n}=(2n-1)\pi/(2\left|\mathbf{d}_{k_{c}}(\gamma_{f})\right|)$ are critical times when DQPTs happen ZengYM . Since $\sin^{2}(\left|\mathbf{d}_{k_{c}}(\gamma_{f})\right|t_{c,n})$ always equals 1, it is easy to get $\mathcal{L}_{k_{c}}(t_{c,n})=(\hat{\mathbf{d}}_{k_{c}}(\gamma_{i})\cdot\mathbf{\hat{d}}_{k_{c}}(\gamma_{f}))^{2}$. Evidently, $\mathcal{L}_{k_{c}}(t_{c,n})=0$ is fulfilled if and only if $\hat{\mathbf{d}}_{k_{c}}(\gamma_{i})\cdot\mathbf{\hat{d}}_{k_{c}}(\gamma_{f})=0$; i.e., $\hat{\mathbf{d}}_{k_{c}}(\gamma_{i})$ and $\mathbf{\hat{d}}_{k_{c}}(\gamma_{f})$ are perpendicular on the Bloch sphere Dora2015PRB . Whereas what we care about is whether such a $k_{c}$ exists for various values of $\gamma_{i}$ and $\gamma_{f}$, we take no account of values of critical times $t_{c,n}$ when $\mathcal{L}_{k_{c}}$ equals zero, and investigate properties of $\bar{\mathcal{L}}_{k}=\mathcal{L}_{k}(\pi/[2\left|\mathbf{d}_{k}(\gamma_{f})\right|])$ which is the minimum of $\mathcal{L}_{k}(t)$ in time instead of $\mathcal{L}_{k}(t)$. According to Eq. (7), it is easy to get
$\bar{\mathcal{L}}_{k}=(\hat{\mathbf{d}}_{k}(\gamma_{i})\cdot\mathbf{\hat{d}}_{k}(\gamma_{f}))^{2},$ (8)
which is time independent and equals zero only for the $k_{c}$-mode, i.e., $\bar{\mathcal{L}}_{k_{c}}=0$.
@@
where $|\psi_{i,k}^{0}\rangle$ and $|\psi_{f,k}^{0}\rangle$ are ground states of prequench Hamiltonian $H_{i,k}$ and postquench Hamiltonian $H_{f,k}$, respectively. Comparing Eq. (9) with Eq. (1), the difference is that the two parameters of the quench fidelity are fixed to the prequench parameter and postquench parameter, respectively. Then we have the $k$-mode of the quench fidelity:
@@
For $\gamma_{i}$ and $\gamma_{f}$ being in different phases, $\mathcal{F}^{\text{q}}$ has at least one $k_{0}$-mode that fulfills $\mathcal{F}^{\text{q}}_{k_{0}}=0$; thus there exists at least one $k_{0}$-mode that fulfills $\hat{\mathbf{d}}_{k_{0}}(\gamma_{i})\cdot\mathbf{\hat{d}}_{k_{0}}(\gamma_{f})=-1$; i.e., $\hat{\mathbf{d}}_{k_{0}}(\gamma_{i})$ and $\hat{\mathbf{d}}_{k_{0}}(\gamma_{f})$ are antiparallel on the Bloch sphere. As a consequence, the quench fidelity $\mathcal{F}^{\text{q}}$ equals to zero indicates the occurrence of an EQPT; i.e., $\gamma_{i}$ and $\gamma_{f}$ lie in different phases. If $\gamma_{i}$ and $\gamma_{f}$ are within the same phase, i.e., no EQPT occurs, then $\mathcal{F}^{\text{q}}$ is nonzero.
Comparing Eq. (10) with Eq. (8), it is straightforward to get
$\bar{\mathcal{L}}_{k}=\left[2(\mathcal{F}^{\text{q}}_{k})^{2}-1\right]^{2}.$ (11)
Equation (11) implies that the $k_{c}$-mode of the LE fulfilling $\bar{\mathcal{L}}_{k_{c}}=0$ corresponds to the $k_{c}$-mode of $\mathcal{F}^{\text{q}}$ fulfilling $\mathcal{F}^{\text{q}}_{k_{c}}=\sqrt{2}/2$. Therefore, a necessary and sufficient condition for the occurrence of DQPTs is that the quench fidelity of the prequench system and the postquench system has at least a $k_{c}$-mode which satisfies $\mathcal{F}^{\text{q}}_{k_{c}}=\sqrt{2}/2$. For a two-band system, if $d_{0,k_{c}}(\gamma_{f})=0$, then
$H_{k_{c}}(\gamma_{f})\left|\psi_{k_{c}}^{0}(\gamma_{i})\right\rangle\propto\left|\psi_{k_{c}}^{1}(\gamma_{i})\right\rangle$ (12)
can be derived from $\mathcal{F}_{k_{c}}=\sqrt{2}/2$, where $\left|\psi_{k_{c}}^{1}(\gamma_{i})\right\rangle$ is the first excited state of $H_{k_{c}}(\gamma_{i})$. This implies $\left\langle\psi_{k_{c}}^{0}(\gamma_{i})\right|H_{k_{c}}(\gamma_{f})\left|\psi_{k_{c}}^{0}(\gamma_{i})\right\rangle=0$. And further $\left\langle\psi_{k_{c}}^{0}(\gamma_{i})\right|H_{k_{c}}^{2}(\gamma_{f})\left|\psi_{k_{c}}^{0}(\gamma_{i})\right\rangle=\left|\mathbf{d}_{k_{c}}(\gamma_{f})\right|^{2}$ can be derived. If $d_{0,k_{c}}(\gamma_{i})$ is also equal to 0, then $H_{k_{c}}(\gamma_{i})$ anticommutes with $H_{k_{c}}(\gamma_{f})$, i.e.,
$\{H_{k_{c}}(\gamma_{i}),H_{k_{c}}(\gamma_{f})\}=0.$ (13)
The detailed derivation of these formulas is given in the Appendix A.